OpenStudy (anonymous):

f(x)=6sin(6 pi x) deterime slope.

OpenStudy (anonymous):

Looking for help here, I know min and max are 6, -6 but how do you determine other points to define slope. I came to -12/(-2pi/12) and reduced -6/-(pi/6) but not working

OpenStudy (tkhunny):

Determine the slope WHERE? Everwhere?

OpenStudy (anonymous):

On the graph of F(x)=6sin(6pix), points P and Q are at consecutive lowest and highest points with P occuring before Q. Find the slope of the line which passes through P and Q

OpenStudy (anonymous):

this is a periodic function, with period \(2\pi/6\pi=1/3\) thhis means that the poinys P and Q are separated by distance \(1/2 \) of the period. So it means 1/6. Now, as you already determined the function max and min values are 6 and -6. Remmember noe the slope formula. slope \(=(y_2-y_1)/(x_2-x_1)\) in your case \(y_2-y_1= -6 -6=-12\) and \(x_2-x_1=1/6\) so....

OpenStudy (anonymous):

sry, you sa P is min and Q max, so \(y_2=Q\) and \(y_1=P\) so \(y_2-y_1=6-(-6)=12\). My bad

OpenStudy (anonymous):

Thank you I have been stewing over that for a few hours ! I was almost there but lacked that little step!